OpenStudy (anonymous):

start from 110 feet away, a person on a bicycle rides towards a checkpoint and then passes it. the rider is travelling at a constant rate of 30 feet per second. the distance between the bicycle and the checkpoint is given by the equation d=110-30t . at what times is the bike 40 feet away from the checkpoint? a. 2.3 sec and 5 sec b. 1.6 sec and 5 sec c. 2.3 sec and 4.6 sec d.1.6 sec and 4.6 sec

OpenStudy (perl):

solve 40 = 110 - 30*t and -40 = 110 - 30*t

OpenStudy (perl):

so i get t = 5 , 2.3

OpenStudy (perl):

you can be 40 feet to the right or left of the origin. thats why you have two equations

OpenStudy (anonymous):

ok got it, so A?